# Equidistribution of polynomially bounded o-minimal curves in homogeneous spaces

### v1

@@ 4. Non-escape of mass
Let $X\cup\{\infty\}$ be the one-point compactification of
@@
Let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a continuous unbounded curve definable in a polynomially bounded o-minimal structure, and for $x\in X$ recall the measures $\mu_{T,\varphi,x}$ defined in (1.1).
By the Banach-Alaoglu theorem, any weak-star limit $\mu$ as $T\to\infty$ of the measures $\mu_{T,\varphi,x}$ is a probability measure on $X\cup\{\infty\}$. First will show the following. Then we will extend the result to the case of finite volume quotient spaces of Lie groups.
@@
If $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ is a continuous unbounded non-contracting curve definable in a polynomially bounded o-minimal structure. Then, there exists $T_{0}>0$ such that given a compact set $F\subset\text{SL}(n,\mathbb{R})$ and $\epsilon>0$, there exists a compact set $K\subset\text{SL}(n,\mathbb{R})/\text{SL}(n,\mathbb{Z})$ such that for any $g\in F$ and $T\geq T_{0}$, we have
@@
We give some preliminaries before the proof. In the discussion that follows, we will refer to rank-$k$ discrete subgroups of $\mathbb{R}^{n}$ as $k$-lattices. For a $k$-lattice $\Lambda\subset\mathbb{Z}^{n}$, let $\|\Lambda\|$ be the volume of a fundamental domain $F\leq\text{Span}_{\mathbb{R}}\{\Lambda\}$ of $\Lambda$ with respect to the usual measure on the subspace $\text{Span}_{\mathbb{R}}\{\Lambda\}\leq\mathbb{R}^{n}$ (which is obtained by restricting the usual Euclidean inner product). Recall that if $\{\textbf{v}_{1},...,\textbf{v}_{k}\}$ forms a $\mathbb{Z}$-basis for $\Lambda$, then
$\|\Lambda\|=\|\textbf{v}_{1}\wedge\cdots\wedge\textbf{v}_{k}\|,$
where the norm is the standard Euclidean norm defined through the inner product for which the pure wedges of $k$-positively oriented tuples of the canonical basis vectors $e_{i},i=1,...,n$ are orthogonal and have norm equal to one. Consider the representation of $\text{SL}(n,\mathbb{R})$ on $\bigwedge^{k}\mathbb{R}^{n}$ defined by
(4.1) $g.(\textbf{v}_{1}\wedge\textbf{v}_{2}...\wedge\textbf{v}_{k}):=g\textbf{v}_{1}\wedge g\textbf{v}_{2}\wedge\cdots\wedge g\textbf{v}_{k},~{}g\in\text{SL}(n,\mathbb{R}).$
For a fixed $k<n$, the action of $\text{SL}(n,\mathbb{R})$ is transitive on the space of $k$-lattices of rank $k<n$. We denote $\mathbb{Z}^{k}:=\text{Span}_{\mathbb{Z}}\{e_{1},...,e_{k}\}$, and observe that for a unimodular $n$-lattice $L=g\mathbb{Z}^{n}\leq\mathbb{R}^{n}$, where $g\in\text{SL}(n,\mathbb{R})$, we have that
$g\gamma\mathbb{Z}^{k},\gamma\in\text{SL}(n,\mathbb{Z}),$
is the collection of primitive $k$-sublattices of $L$. The following powerful theorem on quantitative non-divergence due to Kleinbock [Kleinbockclay] will be needed.
@@
Suppose an interval $B\subset\mathbb{R}$, $C,\alpha>0$, $10<\rho<$ 1 and a continuous map $h:B\to\text{SL}(n,\mathbb{R})$ are given. Assume that for any $\gamma\in\text{SL}(n,\mathbb{Z})$, and $1\leq k<n$ we have
@@
    the function $x\to\|h(x)\gamma.e_{1}\wedge\cdots\wedge e_{k}\|$ is $(C,\alpha)$-good on $B$, and
@@
    $\sup_{x\in B}\|h(x)\gamma.e_{1}\wedge\cdots\wedge e_{k}\|\geq\rho^{k}$.
Then, for any $\epsilon<\rho$,
@@
$\mathcal{B}_{\delta}:=\{L\in\mathcal{L}_{n}:(L\smallsetminus 0)\cap B_{\delta}(0)\neq\emptyset\}.$
Here $B_{\delta}(0)\subseteq\mathbb{R}^{n}$ is the ball of radius $\delta$ centered at the origin. By Mahler’s Criterion ([BM00], Theorem 3,2), $\mathcal{L}_{n}\smallsetminus\mathcal{B}_{\delta}$ is compact for all $\delta>0$ and thus $\mathcal{B}_{\delta}$ is a neighborhood of $\infty$. Fix $g\mathbb{Z}^{n}\in\mathcal{L}_{n}$ for $g\in\text{SL}(n,\mathbb{R})$, and consider the measures
@@
This will be concluded by Theorem 4.2 as follows. We first verify the conditions. For $k\in\{1,2,..,n-1\}$, we denote by $\theta_{k}:\text{SL}(n,\mathbb{R})\to\bigwedge^{k}\mathbb{R}^{n}$ the representation (4.1). By Proposition 1.15, there exists $T_{0},C,\alpha>0$ such that for any fixed $k\in\{1,2,..,n-1\}$, and $\gamma\in\text{SL}(n,\mathbb{Z})$ it holds that
(4.5) $\Theta(t)=\|\theta_{k}(\varphi(t)g\gamma)(e_{1}\wedge\cdots\wedge e_{k})\|=\|\varphi(t)g\gamma.(e_{1}\wedge\cdots\wedge e_{k})\|$
is $(C,\alpha)$-good in $[T_{0},\infty)$. Let $F\subset\text{SL}(n,\mathbb{R})$ be a given compact subset. Then
$\rho:=(\inf\{\|\varphi(T_{0})g\gamma.(e_{1}\wedge\cdots\wedge e_{k})\|:g\in F,\,\gamma\in\text{SL}(n,\mathbb{Z})\})^{\frac{1}{k}}>0.$
Thus, by Theorem 4.2, for any $g\in F$,
(4.6) |$|\{t\in[T_{0},T]:\lambda_{1}(\varphi(t)g\mathbb{Z}^{n})\leq\delta\}|\leq C\left(\frac{\delta}{\rho}\right)^{\alpha}(T-T_{0}),$
@@ 4.1. Non-divergence for homogeneous space of Lie groups
In the setting of Lie groups, one obtains the following result.
@@
Let $G$ be a Lie subgroup of $\text{SL}(n,\mathbb{R})$ and let $\Gamma$ be a lattice in $G$. Suppose that $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ is a non-contracting curve definable in a polynomially bounded o-minimal structure $\mathscr{S}$. Assume further that $\varphi([0,\infty))\subset G_{u}$. Then, there exists a $T_{0}>0$ such that given $\epsilon>0$ and a compact set $F\subset G$, there exists a compact set $K\subset G/\Gamma$ such that for all $g\in F$ and all $T>T_{0}$,
@@
We note that the assumption that $\varphi([0,\infty))\subset G_{u}$ is non restrictive for our purposes in view of the following. In Proposition 5.13, we show that for a curve contained in a Lie subgroup $G\subseteq\text{SL}(n,\mathbb{R})$ which is non-contracing and definable in a polynomially bounded o-minimal structure, it holds that its hull is contained in $G_{u}$. Thus, by multiplying the original curve with a correcting curve, the assumption $\varphi([0,\infty))\subset G_{u}$ is satisfied.
@@
We follow the arguments as in [Dani1993LimitDO, Theorem 6.1] to reduce the problem to the case when $G$ is a semisimple group and $\Gamma$ is an irreducible lattice in $G$.
We note that if $G/\Gamma$ is compact, there is nothing to prove. So, we now assume that $G/\Gamma$ is not compact. Let $M$ be the smallest closed normal subgroup of $G$ such that $\bar{G}=G/M$ is semisimple with the trivial center and no compact factors. Then $M/(M\cap\Gamma)$ is compact, see [RA72, Theorem 8.24] and [Shah1996LimitDO, Proof of Theorem 2.2]. Let $q:G\to\bar{G}$ denote the quotient homomorphism. Then $q(\Gamma)$ is a lattice in $\bar{G}$, and the natural quotient map $\bar{q}:G/\Gamma\to\bar{G}/q(\Gamma)$ is proper. We note that if $\tilde{G}$ and $\tilde{M}$ denote the Zariski closures of $G$ and $M$ in $\text{SL}(n,\mathbb{R})$, respectively, then $\bar{G}=G/M=\tilde{G}^{0}/\tilde{M}^{0}$. Therefore, $q|_{G_{u}}:G_{u}\to\bar{G}$ is a rational homomorphism of real algebraic groups. Thus $q\circ\varphi$ is a curve definable in $\mathscr{S}$. By Lemma 2.15, $q\circ\varphi:[0,\infty)\to\bar{G}$ is a non-contracting. Since $\bar{q}$ is proper, proving the theorem for $\bar{G}$ in place of $G$ is sufficient.
Therefore, we assume that $G$ is semisimple with the trivial center and no compact factors. Then, by [RA72, Theorem 5.22], there exist closed normal subgroups $G_{1},\ldots,G_{k}$ of $G$ for some $k$ such that $G$ equals the direct product $G_{1}\times\cdots\times G_{k}$, such that for $\Gamma_{i}=G_{i}\cap\Gamma$ is an irreducible lattice in $G_{i}$. Moreover, $\Gamma_{1}\times\cdots\times\Gamma_{k}$ is a normal subgroup of finite index in $\Gamma$. Hence, $G/\Gamma$ is finitely covered by $G_{1}/\Gamma_{1}\times\cdots\times G_{k}/\Gamma_{k}$. Let $q_{i}:G\to G_{i}$ denote the natural factor map, and observe that $q_{i}\circ\varphi$ is a non-contracting curve in $G_{i}$ definable in $\mathscr{S}$ by Lemma 2.15. So, proving the theorem for $G_{i}/\Gamma_{i}$ separately for each $i$ will imply the theorem for $G/\Gamma$. Therefore, without loss of generality, we may assume that $\Gamma$ is an irreducible lattice in $G$, which is semisimple with the trivial center and no compact factors.
First, suppose that the real rank of $G$ is at least $2$. Then, by the arithmeticity theorem due to G. A. Margulis, $\Gamma$ is arithmetic; that is, see [Zimmer:book84, page 3]: There exists an $m\in\mathbb{N}$, an algebraic semisimple group $H\subset\text{SL}(m,\mathbb{R})$ defined over $\mathbb{Q}$ and a surjective algebraic homomorphism
@@
such that $\ker\rho$ is compact and $\rho(H^{0}(\mathbb{Z}))$ is commensurable with $\Gamma$, where $H^{0}(\mathbb{Z})=H^{0}\cap\text{SL}(m,\mathbb{Z})$. Since $G$ has no compact factors, the map $\rho$ restricted to $H_{u}$ is a finite cover of $G$. By Lemma 2.19, we can lift $\varphi$ to a $\mathscr{S}$-definable curve $\psi:[0,\infty)\to H_{u}\subset H^{0}\subset\text{SL}(m,\mathbb{R})$ which is non contracting in $H_{u}$, such that $\rho\circ\psi=\varphi$. By Proposition 2.20, $\psi$ in non-contracting in $\text{SL}(m,\mathbb{R})$. As in Proposition 4.1, let $T_{0}>0$ be such that for any compact set $F_{1}\subset\text{SL}(m,\mathbb{R})$, there exists a compact set $K_{1}\subset Y$ such that for any $T\geq T_{0}$ and $h\in F_{1}$
@@
Choose, $F_{1}=\rho^{-1}(F)\subset H^{0}$. Since $H^{0}/H^{0}(\mathbb{Z})\hookrightarrow Y:=\text{SL}(m,\mathbb{R})/\text{SL}(m,\mathbb{Z})$ is a proper $H^{0}$-equivariant continuous injection, we can pick a compact set $F_{2}\subset H^{0}$ such that $F_{2}\text{SL}(m,\mathbb{Z})\supset K_{1}\cap H^{0}\text{SL}(m,\mathbb{Z})$. Since $\Gamma\cap\rho(H^{0}(\mathbb{Z}))$ contains a subgroup of finite index in $\rho(H^{0}(\mathbb{Z}))$, we can pick a finite set $S\subset H^{0}(\mathbb{Z})$ such that $S\rho^{-1}(\Gamma)\supset H^{0}(\mathbb{Z})$. Now $K:=\pi(\rho(F_{2}S))$ is a compact subset of $G/\Gamma$. Then, for any $t\geq T_{0}$ and $g\in F$, let $h\in\rho^{-1}(g)\subset F_{1}$. Let $t\in[T_{0},T]$ such that $\psi(t)h\text{SL}(m,\mathbb{Z})\in K_{1}$. Then $\psi(t)hH^{0}(\mathbb{Z})\in F_{2}H^{0}(\mathbb{Z})$, and hence $\pi(\varphi(t)g)\in K$. Therefore, for any $g\in F$, and $h\in\rho^{-1}(g)\subset F_{1}$, we have
$\{t\in[T_{0},T]:\varphi(t)g\Gamma\in K\}\supset\{t\in[T_{0},T]:\psi(t)h\text{SL}(m,\mathbb{Z})\in K_{1}\}.$
From this and by (4.8), we obtain (4.7). This completes the proof when the real rank of $G$ is at least $2$.
Now suppose that the real rank of $G$ is $1$. It is straightforward to adapt the proof of [Dani:rank-1, Proposition 1.2] to conclude (4.7). For this purpose, it is enough to use the following property of the map $\varphi$ in place of [Dani:rank-1, Lemmas 2.5 and 2.7]: Let $V=\wedge^{d}\mathfrak{g}$, where $d$ is the dimension of the maximal unipotent subgroup of $G$ and $\mathfrak{g}$ the Lie algebra of $G$. Now, $G$ acts on $V$ via $\wedge^{d}\text{Ad}$. Then there exists $T>0$, $C>0$ and $\alpha>0$ such that for any $g\in G$, the map $t\mapsto\lVert\varphi(t)g.p\rVert$ has the $(C,\alpha)$-good property on $[T_{0},\infty)$ by Proposition 1.15.
@@
If $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a continuous unbounded non-contracting curve definable in a polynomially bounded o-minimal structure. Let $G\leq\text{SL}(n,\mathbb{R})$ be a Lie subgroup, $\Gamma$ be a lattice in $G$ and suppose that $\varphi([0,\infty))\subset G_{u}$. Then any weak-* limit $\mu$ of $\mu_{T,\varphi,x}$ has $\mu(\infty)=0$ for all $x\in G/\Gamma$.

### v2

@@ 4. Non-escape of mass
Let $X^{*}:=X\cup\{\infty\}$ be the one-point compactification of
@@
In what follows, $\mathscr{S}_{\mathrm{pbd}}$ is a fixed polynomially bounded o-minimal structure. Let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, continuous and unbounded curve. For $x\in X$ and $T>0$ let $\mu_{T,\varphi,x}$ be the measure defined in (1.1). By the Banach-Alaoglu theorem, any weak-star limit $\mu$ as $T\to\infty$ of the measures $\mu_{T,\varphi,x}$ is a probability measure on $X^{*}$. We now state and prove the non-divergence property in the setup of SL(n,)/$\text{SL}(n,\mathbb{R})/\text{SL}(n,\mathbb{Z})$. Later in the section, we extend the result to the case of finite volume quotient spaces of Lie groups. In the general case, when the Lie group is of higher rank, the non-divergence result builds on the non-divergence in SL(n,)/$\text{SL}(n,\mathbb{R})/\text{SL}(n,\mathbb{Z})$ by using Margulis’ arithmeticity theorem see [Zimmer:book84, page 3].
@@
Let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, continuous, unbounded, non-contracting curve. Then, there exists $T_{0}>0$ such that given a compact set $F\subseteq\text{SL}(n,\mathbb{R})$ and $\epsilon>0$, there exists a compact set $K\subseteq\text{SL}(n,\mathbb{R})/\text{SL}(n,\mathbb{Z})$ such that for any $g\in F$ and $T\geq T_{0}$, we have
@@
We need the following preliminaries for the proof. In the discussion that follows, we will refer to rank-$k$ discrete subgroups of $\mathbb{R}^{n}$ as $k$-lattices. For a $k$-lattice $\Lambda\subseteq\mathbb{Z}^{n}$, let $\|\Lambda\|$ be the volume of a fundamental domain $F\subseteq\text{Span}_{\mathbb{R}}\{\Lambda\}$ of $\Lambda$ with respect to the usual measure on the subspace $\text{Span}_{\mathbb{R}}\{\Lambda\}\subseteq\mathbb{R}^{n}$ (obtained by restricting the usual Euclidean inner product). Recall that if $\{\textbf{v}_{1},...,\textbf{v}_{k}\}$ forms a $\mathbb{Z}$-basis for $\Lambda$, then
$\|\Lambda\|=\|\textbf{v}_{1}\wedge\cdots\wedge\textbf{v}_{k}\|.$
The norm above is defined by choosing the inner product for which the pure wedges of $k$-positively oriented tuples of the canonical basis vectors $e_{i},i=1,...,n$ are orthonormal. Consider the representation of $\text{SL}(n,\mathbb{R})$ on $\bigwedge^{k}\mathbb{R}^{n}$ defined by
(4.1) $g\cdot(\textbf{v}_{1}\wedge\textbf{v}_{2}...\wedge\textbf{v}_{k}):=g\textbf{v}_{1}\wedge g\textbf{v}_{2}\wedge\cdots\wedge g\textbf{v}_{k},~{}g\in\text{SL}(n,\mathbb{R}).$
For a fixed $k<n$, the action of $\text{SL}(n,\mathbb{R})$ is transitive on the space of $k$-lattices of rank $k<n$. We denote $\mathbb{Z}^{k}:=\text{Span}_{\mathbb{Z}}\{e_{1},...,e_{k}\}$, and we observe that for a unimodular $n$-lattice $L=g\mathbb{Z}^{n}\leq\mathbb{R}^{n}$, where $g\in\text{SL}(n,\mathbb{R})$, the collection of primitive $k$-sublattices of $L$ is given by
$g\gamma\mathbb{Z}^{k},\gamma\in\text{SL}(n,\mathbb{Z}).$
The following powerful theorem on quantitative non-divergence due to Kleinbock [Kleinbockclay] will be needed.
@@
Let $B\subseteq\mathbb{R}$ be an interval, let $C\in\mathbb{R},\alpha>0$, $10<\rho<$ 1 and let $h:B\to\text{SL}(n,\mathbb{R})$ be a continuous map. Assume that for every $\gamma\in\text{SL}(n,\mathbb{Z})$, and for every $1\leq k<n$ both of the following conditions are satisfied
@@
    the function $x\to\|h(x)\gamma\cdot e_{1}\wedge\cdots\wedge e_{k}\|$ is $(C,\alpha)$-good on $B$, and
@@
    $\sup_{x\in B}\|h(x)\gamma\cdot e_{1}\wedge\cdots\wedge e_{k}\|\geq\rho^{k}$.
Then, for every positive number $\rho$ such that $\epsilon<\rho$, we have
@@
$\mathcal{B}_{\delta}:=\{L\in\mathcal{L}_{n}:(L\setminus\{0\})\cap B_{\delta}(0)\neq\emptyset\}.$
Here $B_{\delta}(0)\subseteq\mathbb{R}^{n}$ is the ball of radius $\delta$ centered at the origin. By Mahler’s criterion, $\mathcal{L}_{n}\setminus\mathcal{B}_{\delta}$ is compact for all $\delta>0$ and thus $\mathcal{B}_{\delta}$ is a neighborhood of $\infty$. Fix $g\mathbb{Z}^{n}\in\mathcal{L}_{n}$ for $g\in\text{SL}(n,\mathbb{R})$, and consider the measures
@@
This will be concluded by Theorem 4.2 as follows. We first verify the conditions. For $k\in\{1,2,..,n-1\}$, consider the representation (4.1). By Proposition 1.13, there exists $T_{0},C,\alpha>0$ such that for any fixed $k\in\{1,2,..,n-1\}$, and $\gamma\in\text{SL}(n,\mathbb{Z})$ it holds that
(4.5) $\Theta(t)=\|\varphi(t)g\gamma\cdot(e_{1}\wedge\cdots\wedge e_{k})\|$
is $(C,\alpha)$-good in $[T_{0},\infty)$. Let $F\subseteq\text{SL}(n,\mathbb{R})$ be a given compact subset. Then, the number $\rho$ defined by
$\rho:=(\inf\{\|\varphi(T_{0})g\gamma\cdot(e_{1}\wedge\cdots\wedge e_{k})\|:g\in F,\,\gamma\in\text{SL}(n,\mathbb{Z})\})^{\frac{1}{k}}$
is positive. Thus, by Theorem 4.2, for any $g\in F$,
(4.6) |$|\{t\in[T_{0},T]:\lambda_{1}(\varphi(t)g\mathbb{Z}^{n})\leq\delta\}|\leq C\left(\frac{\delta}{\rho}\right)^{\alpha}(T-T_{0}).$
@@ 4.1. Non-divergence for homogeneous space of Lie groups
For the following statement, we recall that for a Lie group $G$, $G_{u}$ denotes the Lie-subgroup generated by the one-parameter unipotent subgroups of $G$.
@@
Let $G$ be a Lie subgroup of $\text{SL}(n,\mathbb{R})$ and let $\Gamma$ be a lattice in $G$. Suppose that $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ is a $\mathscr{S}_{\mathrm{pbd}}$-definable, non-contracting curve. Assume further that $\varphi([0,\infty))\subseteq G_{u}$. Then, there exists a $T_{0}>0$ such that given $\epsilon>0$ and a compact set $F\subseteq G$, there exists a compact set $K\subseteq G/\Gamma$ such that for all $g\in F$ and all $T>T_{0}$,
@@
We note that the assumption that $\varphi([0,\infty))\subseteq G_{u}$ is non restrictive for our purposes in view of the following. In Proposition 5.13, we show that the hull of a non-contracting curve, definable in a polynomially bounded o-minimal structure and taking values in a Lie subgroup $G\subseteq\text{SL}(n,\mathbb{R})$, is contained in $G_{u}$. Thus, by multiplying the original curve with a correcting curve, the assumption $\varphi([0,\infty))\subseteq G_{u}$ will be satisfied.
@@
We follow the arguments as in [Dani1993LimitDO, Theorem 6.1] to reduce the problem to the case when $G$ is a semisimple group and $\Gamma$ is an irreducible lattice in $G$. We note that if $G/\Gamma$ is compact, there is nothing to prove. So, we now assume that $G/\Gamma$ is not compact.
Let $M$ be the smallest closed normal subgroup of $G$ such that $\bar{G}=G/M$ is semisimple with the trivial center and no compact factors. Then $M/(M\cap\Gamma)$ is compact, see [RA72, Theorem 8.24] and [Shah1996LimitDO, Proof of Theorem 2.2]. Let $q:G\to\bar{G}$ denote the quotient homomorphism. Then $q(\Gamma)$ is a lattice in $\bar{G}$, and the natural quotient map $\bar{q}:G/\Gamma\to\bar{G}/q(\Gamma)$ is proper. We note that if $\tilde{G}$ and $\tilde{M}$ denote the Zariski closures of $G$ and $M$ in $\text{SL}(n,\mathbb{R})$, respectively, then $\bar{G}=G/M=\tilde{G}^{0}/\tilde{M}^{0}$. Therefore, $q|_{G_{u}}:G_{u}\to\bar{G}$ is a rational homomorphism of real algebraic groups. Thus $q\circ\varphi$ is a curve definable in $\mathscr{S}_{\mathrm{pbd}}$. By Lemma 2.17, $q\circ\varphi:[0,\infty)\to\bar{G}$ is a non-contracting. Since $\bar{q}$ is proper, proving the theorem for $\bar{G}$ in place of $G$ is sufficient.
Therefore, we assume that $G$ is semisimple with the trivial center and no compact factors. Then, by [RA72, Theorem 5.22], there exist closed normal subgroups $G_{1},\ldots,G_{k}$ of $G$ for some $k$ such that $G$ equals the direct product $G_{1}\times\cdots\times G_{k}$, such that $\Gamma_{i}:=G_{i}\cap\Gamma$ is an irreducible lattice in $G_{i}$, for all $i$. Moreover, $\Gamma_{1}\times\cdots\times\Gamma_{k}$ is a normal subgroup of finite index in $\Gamma$. Hence, $G/\Gamma$ is finitely covered by $G_{1}/\Gamma_{1}\times\cdots\times G_{k}/\Gamma_{k}$. Let $q_{i}:G\to G_{i}$ denote the natural factor map, and observe that $q_{i}\circ\varphi$ is a non-contracting curve in $G_{i}$ definable in $\mathscr{S}_{\mathrm{pbd}}$ by Lemma 2.17. So, proving the theorem for $G_{i}/\Gamma_{i}$ separately for each $i$ will imply the theorem for $G/\Gamma$. Therefore, without loss of generality, we may assume that $\Gamma$ is an irreducible lattice in $G$, which is semisimple with the trivial center and no compact factors.
First, suppose that the real rank of $G$ is $1$. It is straightforward to adapt the proof of [Dani:rank-1, Proposition 1.2] to conclude (4.7). For this purpose, it is enough to use the following property of the map $\varphi$ in place of [Dani:rank-1, Lemmas 2.5 and 2.7]: Let $V=\wedge^{d}\mathfrak{g}$, where $d$ is the dimension of the maximal unipotent subgroup of $G$ and $\mathfrak{g}$ the Lie algebra of $G$. Now, $G$ acts on $V$ via $\wedge^{d}\text{Ad}$. Then there exists $T>0$, $C>0$ and $\alpha>0$ such that for any $g\in G$, the map $t\mapsto\lVert\varphi(t)g\cdot p\rVert$ has the $(C,\alpha)$-good property on $[T_{0},\infty)$ by Proposition 1.13.
Now we suppose that the real rank of $G$ is at least $2$. Then, by the arithmeticity theorem due to G. A. Margulis, $\Gamma$ is arithmetic; that is, see [Zimmer:book84, page 3]: There exists an $m\in\mathbb{N}$, an algebraic semisimple group $H\subseteq\text{SL}(m,\mathbb{R})$ defined over $\mathbb{Q}$ and a surjective algebraic homomorphism
@@
such that $\ker\rho$ is compact and $\rho(H^{0}(\mathbb{Z}))$ is commensurable with $\Gamma$, where $H^{0}(\mathbb{Z})=H^{0}\cap\text{SL}(m,\mathbb{Z})$. Since $G$ has no compact factors, the map $\rho$ restricted to $H_{u}$ is a finite cover of $G$. By Lemma 2.21, we can lift $\varphi$ to a $\mathscr{S}_{\mathrm{pbd}}$-definable curve $\psi:[0,\infty)\to H_{u}\subseteq H^{0}\subseteq\text{SL}(m,\mathbb{R})$ which is non-contracting in $H_{u}$, such that $\rho\circ\psi=\varphi$. By Proposition 2.22, $\psi$ is non-contracting in $\text{SL}(m,\mathbb{R})$. Denote $Y:=\text{SL}(m,\mathbb{R})/\text{SL}(m,\mathbb{Z})$. As in Proposition 4.1, let $T_{0}>0$ be such that for any compact set $F_{1}\subseteq\text{SL}(m,\mathbb{R})$, there exists a compact set $K_{1}\subseteq Y$ such that for any $T\geq T_{0}$ and $h\in F_{1}$
@@
Therefore, since $H^{0}/H^{0}(\mathbb{Z})\hookrightarrow Y$ is a proper $H^{0}$-equivariant continuous injection, the conclusion of the proposition holds for $H^{0}$ in place of $G$, $H^{0}(\mathbb{Z})$ in place of $\Gamma$, and $\psi$ in place of $\varphi$. And, since $\rho\circ\psi=\varphi$, the conclusion of the proposition holds for $\rho(H^{0}(\mathbb{Z}))$ in place of $\Gamma$, and hence it holds for $\Gamma$ due to the commensurability. ∎
@@
Let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, continuous, unbounded, non-contracting curve. Let $G\subseteq\text{SL}(n,\mathbb{R})$ be a Lie subgroup, $\Gamma$ be a lattice in $G$, and suppose that $\varphi([0,\infty))\subseteq G_{u}$. Then, any weak-* limit $\mu$ of $\mu_{T,\varphi,x}$ as $T\to\infty$ is a probability measure on $G/\Gamma$, for all $x\in G/\Gamma$.